OpenStudy (anonymous):

simplify 10n^7/15n^3 HELP please

OpenStudy (raden):

10/15 = 2/3 and n^7/n^3 = n^(7-3) = n^4 so, it it is 2/3 n^4

OpenStudy (anonymous):

Thank u